OpenStudy (anonymous):

Hello guys! I have h[x]= cube root of x and I need to find absolute extrema. Please help

OpenStudy (mathteacher1729):

Over what interval?

OpenStudy (anonymous):

great thanks, from -1 to 8

OpenStudy (anonymous):

I know that I start with taking the derivative for the first derivative test

OpenStudy (anonymous):

h'[x] = 1/3x^-2/3 right?

OpenStudy (anonymous):

but from there i'm kinda in the dark.

OpenStudy (mathteacher1729):

Find all points where the first derivative is equal to zero. Those are critical points.

OpenStudy (anonymous):

thats what I dont understand I don't know how solve 1/3x^-2/3 for 0 :(

OpenStudy (mathteacher1729):

\[\huge f(x) = x^{1/3}\qquad f'(x) = \frac{1}{3}x^{-2/3}\] When does \(f'(x) = 0\) ? \[\huge 0 =\frac{1}{3}x^{-2/3}\] multiply by 3 on both sides to get \[\huge 0 = x ^{-2/3} \Rightarrow 0 = \frac{1}{x^{2/3}}\] That is never true (can't have 0 in the denominator). So there are no critical points. That means you must test the endpoints next.